OpenStudy (anonymous):

Which of the following expressions does not mean the same as the other three? x- 17 x+ (-17) x- (-17) - 17 +x

OpenStudy (anonymous):

please help me im stuck

OpenStudy (anonymous):

if you add a negative number, that is the same as subtracting, right?

OpenStudy (anonymous):

I still dont get it?

OpenStudy (anonymous):

can you please explian a little

OpenStudy (anonymous):

if you just had numbers... then can you see why 4 - 2 means the same as 4 + (-2) ?

OpenStudy (anonymous):

and that if you just rearranged it, it would be the same as -2 + 4 ?

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

-2 + 4 = 2 right? and 4 - 2 = 2 and 4 + (-2) means the same thing... 4 + (-2) = 4 - 2 = 2 Think of a number line... adding a number moves you to the right, and subtracting moves back to the left, but if you add a negative number, you also move back to the left. Adding a negative is the same as subtracting a positive.